# Privacy Technologies for Financial Intelligence

### v1

@@ 3.1 Hashing Techniques
Hashing has a long history in Computer Science. Hashing was first introduced in 1953 to speed up the searching for information in storage systems [Kno75]. The general idea of hashing, or more specifically a hash function, is to map values that are of arbitrary sizes, such as textual strings that can have different lengths, to fixed sized values such as integer numbers that are in a specific range. A hash function therefore converts input values into hash values or hash codes. These hash values are then used as indexes or pointers into a hash table of fixed size. According to [CZ17], a good hash function needs to: (1) map values to hash values as uniformly as possible across the space of possible hash values, (2) the same input value will always result in the same hash value, (3) the mapping process should be fast, and (4) the size of the hash table for a given set of values can be specified by setting the range of integer hash values.
Hashing has been and continues to be a valuable security mechanism for making data unreadable to the human eye, preventing its interception by malicious individuals, and providing a way to validate its integrity. Next we explain few variations of hashing that are useful in the financial intelligence domain.
Keyed Cryptographic Hashing
A keyed hash function (also known as a hash message authentication code, or HMAC) is an algorithm that uses a cryptographic key and a cryptographic hash function to produce a message authentication code that is keyed and hashed [BCK96]. In general HMAC is used to verify the integrity and authenticity of data by computing a fixed-size hash value based on a secret key and input data.
Unlike traditional cryptographic hash functions that generate a hash value solely based on the input data, keyed hashing incorporates a secret key or salt value into the hashing process, making it resistant to tampering and unauthorised modification. The key serves as a shared secret between the sender and receiver, ensuring that only parties possessing the key can compute or verify the hash value. This provides a strong level of security against malicious tampering or forgery attempts, as any alteration to the input data or hash value would require knowledge of the secret key to generate a valid HMAC. Keyed cryptographic hashing is widely used in various security applications, including message authentication, data integrity verification, and secure communication protocols [KSN+23].
Bloom Filters
A Bloom filter [Blo70] is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It is implemented as a bit array $BF[i]$, where $i\in[1,n]$, consisting of $n$ bits. The Bloom filter uses $k$ independent hash functions $H_{1}(x),\cdots,H_{k}(x)$, each of which maps an element $x$ in a set $S$ of $m$ elements to one of the $n$ bits within the bit array $BF$. To determine if an element $x$ is present in the Bloom filter, one needs to check if all $k$ hash functions map $x$ to bits that are set to 1 in the Bloom filter. Because of hash collisions, false positives can occur in a Bloom filter, but false negatives are impossible.
Figure 5: The Dice similarity calculation between the names ‘peter’ and ‘pete’, converted into bigrams ($q=2$) and then encoded into two Bloom filters $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ of length $l=12$ bits using $k=2$ hash functions. The bit “$1$” shown in italics at position $6$ in $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ is a hash collision, because both ‘pe’ and ‘te’ are hashed to this position.
In the privacy-preserving record linkage domain, Bloom filters are commonly used for matching string values or computing similarities between string pairs in a privacy-preserving manner [SBR09, CRS20]. When matching string values, each string is first converted into a set of q-grams, where each q-gram is typically of length 2 or 3. These q-grams are then encoded into a Bloom filter using $k$ hash functions. To compute similarities between Bloom filters, a set-based similarity function such as Jaccard or Dice can be used. Recent literature has also shown that Bloom filters can be used to match both numerical and textual values in databases [WVKR23]. Figure 5 illustrates an example of using a Bloom filter for data matching. The Privacy-Preserving Record Linkage step in Figure 4 can be solved using variations of Bloom Filters. The hash functions in such Bloom Filters are usually keyed cryptographic hashing functions to protect against dictionary attacks.
3.2 Homomorphic Encryption
Homomorphic encryption (HE) is a class of encryption schemes that allows certain arithmetic operations to be performed directly on the ciphertext [AAUC18]. HE schemes are usually public-key cryptosystems. Let $m_{1}$ and $m_{2}$ be two plaintexts, and let $pk$ and $sk$ be the public key and secret key for the encryption Enc() and decryption Dec() functions, respectively. The correctness of a HE scheme is expressed by the following property. For an operation $\diamond$ in the plaintext space, there is a corresponding operation $\bullet$ in the ciphertext space such that
$\text{Dec}(sk,\text{Enc}(pk,m_{1})\bullet\text{Enc}(pk,m_{2}))=m_{1}\diamond m_{2},$ (1)
where the different symbols indicate that the arithmetic operations, either addition or multiplication, in the plaintext and ciphertext spaces may differ. The term “homomorphic” reflects that the decryption function acts like a homomorphism, preserving the structure of the mathematical object.
Depending on the types of arithmetic operations permitted and the number of operations it can handle before decryption fails, an HE scheme is classified into one of the following types, where the algorithmic complexity of the operations we want to perform on data is measured in terms of the depth of circuits and the arithmetic gates supported.
  • Partially HE (PHE) - Schemes that can evaluate circuits containing only one type of arithmetic gate, that is, either addition or multiplication, for unbounded circuit depth.
  • Somewhat HE (SHE) - Schemes that can evaluate a subset of circuits containing both addition and multiplication gates, where the complexity grows as the circuit depth increases.
  • Fully HE (FHE) - Schemes that can evaluate all circuits containing both addition and multiplication gates for unbounded circuit depth.
The privacy-preserving record linkage and model scoring steps in Figure 4 can be done using PHE, but full-scale risk modelling using both supervised and unsupervised machine learning algorithms would usually require SHE or even FHE. FHE was made possible only in recent years due to the breakthrough described in [Gen09]. Gentry’s work is technically based on lattice-based cryptography, an active research area also seen as a strong candidate for quantum-resistant encryption [A+22]. In addition to the standard HE components, Gentry’s scheme introduces two key concepts: squashing and bootstrapping. Simply put, squashing reduces the complexity of decryption by providing a hint to the evaluation process (which introduces additional security assumptions), while bootstrapping allows the scheme to evaluate its own decryption along with an arithmetic operation. When combined, these techniques enable the scheme to perform unbounded arithmetic operations with controlled decryption errors.
It is worth noting that HE schemes are inherently probabilistic, meaning that the same plaintext can produce a different ciphertext each time it is encrypted. This is to ensure a security notion called semantic security. For a more thorough introduction to HE and its mathematical and computer science foundations, the reader is referred to [Hal17, Pei16, LNP22].
To conclude this section, we provide an example of a partially homomorphic encryption (PHE) scheme: the multiplicative HE scheme known as ElGamal Encryption [ElG85]. This scheme is based on the Diffie-Hellman key exchange protocol [DH76] to securely exchange a masked secret key between Alice and Bob for encryption and decryption.
Public key ElGamal Encryption. Bob is the sender, Alice is the receiver. Alice chooses a cyclic group $G$ with order $q$ and a generator $g$.
Private key: Alice’s secret key is a random integer $x\leftarrow\{1,\dots,q-1\}$. Public key: Alice computes $h=g^{x}\in G$ and publishes the public key $(G,q,g,h)$. Encryption: To encrypt a message $m\in G$, Bob randomly selects $y\leftarrow\{1,\dots,p-1\}$, computes the shared secret $s=h^{y}\in G$, and sends the ciphertext $(c_{1},c_{2})$, where $c_{1}=g^{y}$ and $c_{2}=m\cdot s$. Decryption: To decrypt, Alice computes the shared secret $s=c_{1}^{x}$ and its inverse $s^{-1}\in G$, then recovers the message $m=c_{2}\cdot s^{-1}$.
Since each term in the ElGamal ciphertext can be multiplied separately to produce a multiplication ciphertext, and Alice can recover the correct multiplied message using the above decryption process, it implies that ElGamal is a multiplicative HE scheme. Specifically, if $C=(c_{1},c_{2})\leftarrow\text{Enc}(m)$ and $\tilde{C}=(\tilde{c}_{1},\tilde{c}_{2})\leftarrow\text{Enc}(\tilde{m})$, then decryption of the product $C\cdot\tilde{C}$ yields $m\cdot\tilde{m}$, i.e., Dec$(C\cdot\tilde{C})=m\cdot\tilde{m}$. We will see some applications of the additive homomorphic form of ElGamal encryption in Section 4.2.
3.3 Secure Multiparty Computation
Secure multiparty computation (MPC) is a concept that drives the development of protocols designed to securely and accurately perform computations across multiple parties. Research in MPC originated with Yao’s so-called Millionaires’ problem [Yao82], in which two parties seek to know who is richer but without having to disclose their net wealth to each other or a trusted third party. There are many solutions to the Millionaires’ problem involving different cryptographic primitives. Here we briefly describe an accessible solution from [LT05]. The algorithm first converts two numbers in binary representation $x=x_{n}x_{n-1}\ldots x_{1}$ and $y=y_{n}y_{n-1}\ldots y_{1}$ we want to privately compare to the following two sets
$S_{x}=\{x_{n}x_{n-1}\ldots x_{i+1}1\,|\,x_{i}=0,1\leq i\leq n\}$
$S_{y}=\{y_{n}y_{n-1}\ldots y_{i}\,|\,y_{i}=1,1\leq i\leq n\}$
and then securely compute the cardinality $|S_{x}\cap S_{y}|$. The correctness of the procedure comes from the fact that $x>y$ if and only if $S_{x}\cap S_{y}$ is non-empty. There are various private set intersection cardinality (PSI-CA) protocols for computing $|S_{x}\cap S_{y}|$, utilising different techniques such as homomorphic encryption, hashing, Bloom filters, and oblivious transfer. Here, we describe a PSI-CA protocol that employs hashing and commutative encryption from [AES03].
Informally, an encryption function $f$ is commutative if the order of encryption does not matter for arbitrary encryption keys $e$ and $e^{\prime}$, i.e., $f_{e}\circ f_{e^{\prime}}=f_{e^{\prime}}\circ f_{e}$ where $\circ$ is function composition. An example given in [AES03] is the function
$f_{e}(x)=x^{e}\mod p,$
where the powers commute, yielding
$(x^{e}\bmod p)^{e^{\prime}}\bmod p=x^{ee^{\prime}}\bmod p=(x^{e^{\prime}}\bmod p)^{e}\bmod p.$
The security of this commutative encryption is based on the Decisional Diffie-Hellman hypothesis [Bon98]. Armed with a commutative encryption function, $|S_{x}\cap S_{y}|$ can be computed via double encryption as follows:
$|S_{x}\cap S_{y}|=|f_{k_{y}}(f_{k_{x}}(S_{x}))\cap f_{k_{x}}(f_{k_{y}}(S_{y}))|,$
where $k_{x}$ and $k_{y}$ are the encryption keys for the two parties holding the value $x$ and $y$, respectively.
The above secure number comparison protocol provides a glimpse into the rich literature on MPC [Gol98, CD15, EKR+18]. Interestingly, what makes a distributed computation task an MPC protocol cannot be straightforwardly defined in a mathematical way. The well-accepted way of defining MPC’s security is by simulating an ideal world and a real world, known as the ideal/real simulation paradigm [Lin20]. In the simulation, the two worlds are identical except that there is a trusted third party in the ideal world but not in the real world. To examine the security of a protocol under adversarial attacks, one would compare the outcomes of the protocol in the two worlds to determine if the adversary in the real world is able to gather any additional non-disclosed information, should the same attacks also be successful in the ideal world.
With the ideal/real simulation paradigm in place, the following are important properties of an MPC protocol. Privacy guarantees that an adversary learns nothing about an honest party’s input beyond what is prescribed. Correctness ensures that the protocol always produces the correct result despite being attacked by adversaries. Guaranteed output stops the adversary from halting the protocol hence stopping others from receiving an output. Lastly fairness states that an adversary receives the protocol output if and only if an honest party receives the protocol output.
There are two primary adversarial models: semi-honest and malicious adversaries. In the semi-honest model, an adversary correctly follows the protocol specification, but seeks opportunities to learn the protected privacy from information obtained during the computation process. In contrast, the malicious adversarial behaviour can arbitrarily deviate from the protocol, including aborting the protocol or changing its inputs and outputs.
@@ 3.4 Differential Privacy
Differential privacy (DP) is a technique that aims to enable the analysis of sensitive data while protecting the privacy of individuals whose data is included in the analysis [DR14]. In general, DP is achieved by adding noise in a controlled manner during the analysis process, ensuring that statistical queries do not reveal information about specific individuals. This allows organisations to derive valuable insights from their data while minimising the risk of unauthorised disclosure or misuse of sensitive personal data.
Definition 3.4.1.
A randomised algorithm $\mathcal{A}$ is $\epsilon$-differentially private if for all neighbouring databases $\mathbf{D}$ and $\mathbf{D^{\prime}}$ that differ in at most one record, and for all sets $\mathcal{O}$ outputs, we have
@@
The parameter $\epsilon$ is known as the privacy budget in DP. A smaller value of $\epsilon$ indicates a stronger privacy guarantee, as it reflects a stricter bound on how much the query results can change in response to the inclusion or exclusion of a single record. The Bayesian semantics of differential privacy can be found in [KS14, KM14, YZN24]; in short, DP guarantees that the odds ratio of an attacker’s prior and posterior beliefs (before and after seeing the query output) in whether a specific record is part of the input database is bounded, and that confers every individual with plausible deniability on whether the person’s data is in the dataset.
An example scenario that illustrates Definition 3.4.1 is differencing attacks. Consider a medical database containing records of individuals who participated in a disease screening. An analyst with access only to counting queries can query the database under different conditions to select two distinct subsets of the data. If the difference between these subsets is designed to isolate a specific individual, the analyst could potentially infer whether or not this individual has the disease by examining the difference between the two query results. Differential privacy prevents this by injecting suitably calibrated noise into the results of these queries, thus preventing the analyst from reliably determining the presence or absence of any particular individual based on the differences in query results.
A fundamental DP technique is the Laplace Mechanism [DMNS06, DR14]. Given a real-valued query function $f:\mathbb{N}^{|\chi|}\rightarrow\mathbb{R}^{k}$, the Laplace Mechanism directly adds Laplace noise to the query outputs as
$M_{L}(\mathbf{D},f,\epsilon)=f(\mathbf{D})+(e_{1},\dots,e_{k}),\text{ where }e_{i}\leftarrow\text{Lap}\left(\frac{\Delta f}{\epsilon}\right),$
where Lap() is the Laplace distribution, and
$\Delta f=\max_{\mathbf{D}\sim\mathbf{D}^{\prime}\in\mathbb{N}^{|\chi|}}{||f(\mathbf{D})-f(\mathbf{D}^{\prime})||_{1}}$
is a measure of the sensitivity of the query on two neighbouring databases $\mathbf{D}$ and $\mathbf{D}^{\prime}$. The independent and identically distributed noise is randomly sampled from the Laplace distribution with mean 0 and scale $\Delta f/\epsilon$. The noise’s scale is proportional to the query sensitivity (when $\epsilon$ is fixed) and inversely proportional to the noise budget $\epsilon$ (when $\Delta f$ is fixed). This makes good sense, since a less sensitive query needs less noise to hide its output variation. A smaller $\epsilon$ indicates stronger privacy guarantee hence more noise needs to be added to the outputs.
In the financial domain, where large volumes of sensitive data are collected and analysed for various purposes such as risk assessment, fraud detection, customer segmentation, and personalised marketing, DP serves as a valuable tool for balancing the need for data-driven insights with privacy protection. By incorporating DP into data analytics workflows, financial institutions can adhere to stringent regulatory requirements, such as GDPR and Basel III [^5] See https://gdpr-info.eu and https://www.bis.org/bcbs/basel3.htm, while still leveraging the richness of their datasets for business intelligence and decision-making [Nis21]. Essentially all financial data are graph data and a survey of DP algorithms for graph data can be found in [LPR+23].
3.5 Federated Learning
Federated learning (FL) [MMR+17] is a new learning paradigm that aims to build a joint machine learning model based on data distributed across multiple sites or owned by different participants. During model training, information such as gradients is exchanged between participants, but the raw data are not shared. The intention is that the exchanged information does not reveal any private data belonging to any party.
Figure 6: A federated learning architecture. In order to guarantee the privacy of the data, federated learning only permits all the data owners to exchange the model gradient ($w$) with the aggregation server. During this process, each data owner trains its own model with the local data, then uploads the local model to the aggregation server. After aggregating all the received models, the server returns the new global model updates ($W$) to each data owner.
Figure 6 illustrates a typical FL process. Initially, each data owner receives a generic global model from the aggregation server. After receiving the initial model, each data owner conducts local training on this model with their own data and then uploads the related gradient information (local model updates) to the aggregation server. The aggregation server averages the updates received from data owners into the global model, then updates the global model to replace each user’s local model. This process repeats until the global model achieves the desired performance or the training reaches the maximum iteration number.
FL can be categorised into three types based on the distribution of data: horizontal FL, vertical FL, and federated transfer learning [LSTS20]. Horizontal FL occurs when the user features of the datasets overlap significantly, but the users themselves overlap minimally. Vertical FL occurs when the user features of the datasets overlap minimally, but the users themselves overlap significantly. In cases where both users and user features of the datasets overlap rarely, transfer learning can be used to address the lack of training data [LKX+20].
Privacy protection is a major concern in FL [ARPS20, NDP+21]. Although FL protects private data on each device by exchanging only model gradients with the aggregation server, the model communication during the entire training process can still potentially leak sensitive information to third parties, for example, through reverse engineering of models [ZLD22]. Two techniques can be employed to prevent private data leakage during FL model transmission:
  • Secure aggregation techniques ensure that raw data or local updates remain private during the aggregation process. Methods such as MPC or HE (described in Sections 3.3 and 3.2) can be used to aggregate updates without revealing individual contributions.
  • DP methods add noise to local updates to protect individual privacy. This noise ensures that the contribution of any single device is indistinguishable from the contributions of others. Aggregating noisy updates provides privacy guarantees while still allowing the central model to learn from the collective data.
In federated learning, model scoring and model learning are two distinct phases of the process. Model scoring refers to the evaluation or assessment of the performance of a machine learning model using a set of evaluation metrics. Common evaluation metrics for model scoring include accuracy, precision, recall, F1-score, and area under the ROC curve (AUC). Both federated model learning and federated model scoring are important steps in the future-state financial intelligence proposed in Figure 4.
@@ 4 Privacy-Preserving Information Sharing
In this section, we describe several real-world financial intelligence scenarios wherein information-sharing between different parties is crucial to support effective detection of financial crimes. In each scenario, we discuss the privacy risks during multiparty collaborations and examine the feasibility of applying various privacy-preserving techniques to address these privacy risks.
@@ 4.1 Suspicious Persons List
Suppose an FIU possesses a moderate-sized list of suspicious individuals (or, more generally, commercial entities) obtained, say, from its own investigations and law-enforcement partner agencies and the FIU wishes to acquire additional information about those suspicious individuals, including where they conduct their financial activities and whether they are known to other FIUs. This can be done if the FIU can safely distribute the Suspicious Person List (SPL) to other FIUs and financial institutions to verify the association of individuals on the list with specific financial entities before engaging in further targeted information requests or exchanges with those financial entities.
@@
The problem described above of identifying common individuals between the SPL and each participating financial institution’s customer database can be formulated as a private set intersection (PSI) problem. In a two-party computation protocol, the goal of PSI is to identify the common elements that appear in both parties’ sets and nothing else.
We now describe a simple semi-honest PSI protocol from [FNP04] that utilises polynomial evaluation and homomorphic encryption to ensure the protection of unmatched individuals in the SPL. For detailed security and efficiency analyses, readers are encouraged to refer to [FNP04]. For simplicity, it is assumed that each individual can be uniquely identified by an identification (ID) number, which can be derived from a national ID (in some countries) or a combination of personal information. Here are the steps of the protocol.
@@
    which is uniquely defined according to the Fundamental Theorem of Algebra. The FIU then encrypts the coefficients of $f(x)$ using HE and send the encrypted coefficients and the public key to the other parties:
@@
    Once the matching outputs are shared, FIU decrypts the outputs and find those elements that also appear in its suspicious person list.
@@
The FIU.net operated jointly by FIUs in the European Union has offered privacy-preserving sharing of suspicious entities for over a decade now, using the Ma3tch algorithm [BM13]. A recent legal review of FIU.net can be found in [Mou20]. In Australia, the Australian Financial Crime Exchange [^6] http://afcx.com.au is a platform for financial institutions to share lists of entities of interest, from a financial crime and cyber crime perspective, with each other and, selectively, with the Australian Taxation Office.
There are now many fast PSI algorithms that can operate at large scale, including [CLR17, RT21, CM20]. A key issue with PSI is that an adversary can use repeated PSI computations to completely recover an organisation’s entire dataset. Proposals like [PLN23] seek to address this issue by introducing DP protection into PSI computations.

### v2

@@ 3.3 Hashing Techniques
In cross-institutional data collaboration, a fundamental but often overlooked problem is how to preserve matchability between entities or relationships while avoiding direct exposure of sensitive identifiers. At this level, hashing provides a lightweight yet practical solution. Hashing works by applying a one-way mapping to original identifiers such as customer IDs, transforming them into fixed length representations. For example,
$\text{Hash: Hello}\rightarrow\text{2CF24DBA5FB0A30E26E83B2AC5B9E}$
While the original identifiers cannot be recovered, consistent hashing allows entities or relationships across different datasets to be matched. In other words, hashing does not hide whether two records refer to the same entity, it hides who that entity is.
From a functional standpoint, hashing is not a joint computation mechanism and does not attempt to provide comprehensive protocol level privacy guarantees. Its primary value lies in reducing direct exposure of sensitive identifiers during data exchange and comparison, thereby offering a baseline level of privacy protection. This makes hashing widely used in practice for identity alignment, relationship linking, and preliminary data filtering.
In financial intelligence contexts, hashing is commonly applied as a preprocessing step for customer or account identifiers. This allows financial institutions and regulators or investigators to assess potential overlaps or business relationships without sharing plaintext identity information. While such approaches do not eliminate all privacy risks, they can significantly reduce unnecessary data exposure when computation goals are well defined and trust boundaries are controlled.
As suggested by Chi et al. [17], properties such as uniform distribution of hash outputs, consistency of mapping, computational efficiency, and appropriate handling of input and output domains are important when selecting a hash function for practical applications. More broadly, the effectiveness of hashing depends on several underlying assumptions, including the size of the input space, the consistency of hashing rules across participants, and resilience against auxiliary information attacks. In higher-risk or more complex multiparty collaboration settings, hashing alone may be insufficient; it is therefore typically combined with stronger privacy-preserving mechanisms rather than deployed as a standalone solution.
3.4 Federated Learning
In some collaborative settings, privacy risks arise not from a single joint computation, but from long term, iterative model training processes. In such cases, participants seek to collaboratively build and update models over time while keeping their data within local domains, rather than jointly evaluating a fixed function. Federated learning (FL) is designed for precisely this type of collaboration.
The core idea of federated learning is that data remains local to each participant, model training is performed independently on local datasets, and only model updates are shared with a coordinating entity for aggregation into a global model as shown in Figure 4. From a collaboration perspective, FL does not require sharing raw data or encrypting an entire joint computation. Instead, it follows a local computation and global aggregation paradigm.
Figure 4: An example federated learning (FL) architecture with $n$ data owners. In the setting, data owners only exchange the model gradient ($w$) with the aggregation server. In each training iteration, each data owner $i$ trains its own model with their local data and send the corresponding local model updates $w_{i}$ to the aggregation server. After aggregating all the received local model updates, the aggregations server returns the new global model updates ($W$) to each data owner.
This leads to a fundamental structural difference between FL and MPC. While MPC focuses on securely evaluating a predefined function, FL is concerned with the continuous evolution of a model over time. In large scale, long running, and frequently updated training tasks, attempting to replicate FL using MPC would be impractical due to excessive system complexity, communication overhead, and reduced training efficiency.
It should be noted, however, that although FL avoids direct data sharing, privacy risks do not disappear entirely. Model parameters or gradient updates may, under certain conditions, leak statistical properties or even individual level information about local datasets [3, 46, 66]. Consequently, FL is often combined with additional privacy technologies, such as homomorphic encryption or differential privacy, to mitigate information leakage during the training process.
Within financial intelligence, FL is particularly well suited to scenarios where multiple institutions wish to jointly train models for risk detection, anomaly identification, or behavioural analysis without centralising sensitive data. Participants retain control over their data while benefiting from improved model performance through collaboration. In such settings, the key challenge is not how to securely compute a single result, but how to balance model utility and privacy risk over sustained collaboration, a challenge for which FL offers distinct advantages over MPC.
@@ 3.5 Differential Privacy
Across the privacy technologies discussed so far, the primary focus has been on protecting input data and ensuring trustworthy computation. However, in many real world applications, the results themselves can still leak sensitive information, even when inputs and computation processes are fully protected. Differential privacy (DP) addresses this class of risks directly.
The core principle of DP is to introduce carefully calibrated randomness into statistical outputs or model results, ensuring that the presence or absence of any single individual’s data does not have a distinguishable impact on the output.
Definition 1
A randomised algorithm $\mathcal{A}$ is $\epsilon$-differentially private if for all databases $\mathbf{D}$ and $\mathbf{D^{\prime}}$ that differ in at most one record, and for all sets $\mathcal{O}$ outputs, we have
@@
Rather than hiding raw data or computation steps, DP explicitly constrains how much information can be inferred from released results.
This positioning fundamentally distinguishes DP from MPC. Even in a fully trusted joint computation environment, repeated queries or downstream inference based on exact outputs can gradually expose individual level information. DP mitigates this risk by limiting the sensitivity of outputs to individual data points, something that MPC, HE, or FL cannot achieve on their own.
In financial intelligence applications, DP is commonly used when publishing aggregate statistics, analytical reports, or model outputs. For example, when sharing risk distributions, trend analyses, or group level behavioural insights, DP mechanisms can preserve the overall analytical value while substantially reducing the risk of re-identifying individual customers or accounts. In these contexts, the primary privacy concern is not who performed the computation, but what can be inferred once the results are released or reused [47, 38].
It is important to emphasise that differential privacy is not a zero cost solution. Introducing randomness to limit information leakage necessarily involves a trade-off between privacy guarantees and result accuracy. As such, DP is best understood as an output level privacy governance mechanism, complementing rather than replacing input level and computation level privacy technologies.
@@ 4 Privacy-Preserving Information Sharing
In this section, we revisit the three computational problems listed at the end of § 2 and further contextualise them using real-world financial intelligence scenarios wherein information-sharing between different parties is crucial to support effective detection of financial crimes. In each scenario, we discuss the privacy risks during multiparty collaborations and examine the feasibility of applying various privacy-preserving techniques to address these privacy risks.
@@ 4.1 Suspicious Persons List
Suppose an FIU possesses a moderate-sized list of suspicious individuals (or, more generally, commercial entities) obtained, say, from its own investigations and law-enforcement agencies and the FIU wishes to acquire additional information about those suspicious individuals, including where they conduct their financial activities and whether they are known to other FIUs. This can be done if the FIU can safely distribute the Suspicious Person List (SPL) to other FIUs and financial institutions to verify the association of individuals on the list with specific financial entities before engaging in further targeted information exchanges with those financial entities.
@@
The problem described above of identifying common individuals between the SPL and each participating financial institution’s customer database can be formulated as a private set intersection (PSI) problem. The goal of PSI is to identify the common elements that appear in the parties’ sets and nothing else.
We now describe a simple semi-honest PSI protocol from [26] that utilises polynomial evaluation and homomorphic encryption to ensure the protection of unmatched individuals in the SPL. For detailed security and efficiency analyses, readers are encouraged to refer to [26]. For simplicity, it is assumed that each individual can be uniquely identified by an identification (ID) number, which can be derived from a national ID (in some countries) or a combination of personal information. Here are the steps of the protocol.
@@
    which is uniquely defined according to the Fundamental Theorem of Algebra. The FIU then encrypts the coefficients of $f(x)$ using HE and sends the encrypted coefficients and the public key to the other parties:
@@
    Once the matching outputs are shared, FIU decrypts the outputs and finds those elements that also appear in its suspicious person list.
@@
The FIU.net operated jointly by FIUs in the European Union has offered privacy-preserving sharing of suspicious entities for over a decade now, using the Ma3tch algorithm [6]. A recent legal review of FIU.net can be found in [45]. In Australia, the Australian Financial Crime Exchange (http://afcx.com.au) is a platform for financial institutions to share lists of entities of interest, from a financial crime and cyber crime perspective, with each other and, selectively, with the Australian Taxation Office.
There are now many fast PSI algorithms that can operate at large scale, including [16, 53, 15]. A key issue with PSI is that an adversary can use repeated PSI computations to completely recover an organisation’s entire dataset. Proposals like [50] seek to address this issue by introducing DP protection into PSI computations.